### v2

# Topologically Driven Spin-Orbit Torque in Dirac Matter

Abstract
We unveil novel spin-orbit torque mechanisms driven by topological edge states in magnetic graphene-based devices. Within the energy gap, a damping-like torque plateau emerges within the quantum anomalous Hall phase upon breaking particle-hole symmetry, while for energies at the spin-split Dirac points located within the bands, a large damping-like torque develops as a result of a vanishing Fermi contour. Such torques are tunable by the degree of spin-pseudospin entanglement dictated by proximity-induced spin-orbit coupling terms.
The spin-orbit torque (SOT) mechanism represents a novel method to electrically manipulate a magnetic state, providing power-efficiency and miniaturization prospects beyond traditional multi-ferromagnet setups [1]. While the traditional spin-transfer torque mechanism relies on the generation of spin-polarized currents using ferromagnets [2], the SOT mechanism leverages SOC within metals to generate a current-driven spin source via charge-to-spin conversion [3, 4]. The non-equilibrium spin source $\bm{S}$ couples to the magnetic state, oriented along $\hat{\bm{m}}$, exerting a torque $\bm{T}\propto\hat{\bm{m}}\times\bm{S}$, driving the magnetization dynamics to eventually flip the magnetic state. The conventional SOT phenomenon focuses on heavy metal/ferromagnet bilayers, considering two contributions: the damping-like (DL) torque which stems from a spin current injected from the bulk metal to the ferromagnet via the spin Hall effect, and the field-like (FL) torque generated by an interfacial spin density via the Rashba-Edelstein effect (REE).
Research in alternative systems, such as van der Waals materials and single-layered devices, has revealed that the origin of SOTs could be actually more complex [5, 6]. Despite the suppression of the perpendicular spin Hall effect, DL torques have been measured in single-layered devices [7, 8, 9] and two-dimensional materials [10, 11, 12] and were originally explained by intrinsic Berry curvature effects [7], while the role of non-equilibrium spin textures has recently been revealed [13]. The development of atomically smooth interfaces demonstrates the fundamental role of crystal symmetries for generating unconventional torque contributions, eventually achieving field-free perpendicular magnetization switching [14, 15, 16]; although the spin or orbital nature of such torque remains debated [17]. Furthermore, the interplay between spin and other forms of angular momentum has been shown to enable novel SOT effects [13, 18], where the band structure topology might also play a role [19]. In graphene, spin-orbit coupling (SOC) is mediated by the pseudospin, which corresponds to an additional angular momentum degree of freedom stemming from the lattice [20], and spin-pseudospin correlations have been argued to play a central role in spin transport and charge-to-spin conversion [21, 22, 23]. Indeed, it has been shown that low-energy spin relaxation in graphene is driven by spin-pseudospin locking [21, 23], while the REE is limited by the quenching of the spin texture due to spin-pseudospin entanglement [22]. More recently, spin-pseudospin entanglement was also found to drive SOT beyond the semi-classical spin dynamics for Rashba-like SOC fields [13].
In this Letter we unveil a novel SOT of topological origin generated by spin-pseudospin entanglement in graphene-based heterostructures where magnetic exchange coupling coexists with Rashba and Kane-Mele SOC fields. We first demonstrate the manipulation of spin-pseudospin entanglement and its potential for enhancing charge-to-spin conversion by tuning the interplay between different forms of SOC. By these means, charge-to-spin conversion with maximal efficiency is achieved via the REE in non-magnetic graphene-based heterostructures. Upon including a magnetic exchange coupling we find novel SOT mechanisms which are enabled by tuning spin-pseudospin entanglement at band inversions. Remarkably, we observe a DL torque plateau within the band gap emerging in the insulating topological phase of the quantum anomalous Hall effect, which is activated upon breaking particle-hole symmetry. Additionally, at the spin-split Dirac points a strong DL torque is generated from a vanishing Fermi contour. We trace back the origin of these torques to band structure properties to understand their microscopic origin. Our results unveil novel SOT mechanisms driven by topological phases, as well as manifesting the central role of correlations between spin and other forms of angular momentum for developing optimal spintronic platforms.
We consider a minimal graphene model comprising both Rashba and Kane-Mele SOC with an out-of-plane magnetization. The Hamiltonian is given by
[…]
with $v\sim 10^{6}\,\text{m/s}$ the velocity of massless Dirac electrons in graphene, $\lambda_{\text{R}}$ and $\lambda_{\text{KM}}$ the Rashba and Kane-Mele SOC parameters respectively, and $J_{\text{ex}}$ the exchange splitting along the magnetization direction $\hat{\bm{z}}$. The spin and pseudospin degrees of freedom are respectively represented by the Pauli vectors $\bm{s}$ and $\bm{\sigma}$, while $\tau=\pm$ represents the valley index, where we will focus our analysis on the $\tau=+$ subspace without loss of generality [24]. While SOC in bare graphene is negligible, such SOC fields can be induced by proximity effects in graphene/topological insulator [25, 26, 27, 28, 29] and graphene/transition metal dichalcogenide heterostructures [30, 31, 32, 33, 34, 35]. On the other hand, the magnetization term is obtained by interfacing graphene with an insulating ferromagnet [36, 37, 38, 39, 40, 41].
The exchange splitting is usually the dominant energy scale, splitting the energy dispersion into spin majority and spin minority Dirac cones centered about energy $\mp J_{\text{ex}}/2$. Rashba SOC opens a gap at the crossing between both cones and induces a helical component to the spin texture, as shown in Fig. 1-(a).
The non-equilibrium spin density emerging from the REE, which is a Fermi-surface phenomenon, is computed using a Boltzmann approach as
$\bm{S}_{\text{surf}}(\varepsilon_{\text{F}})=\sum_{n}\int\frac{\text{d}^{2}\bm{k}}{(2\pi)^{2}}\,\delta\mathit{f}_{n,\bm{k}}\bm{s}_{n,\bm{k}}\text{ ,}$ (2)
[…]
$\bm{S}_{\text{sea}}(\varepsilon_{\text{F}})=\sum_{n}\int\frac{\text{d}^{2}\bm{k}}{(2\pi)^{2}}\,\mathit{f}_{n,\bm{k}}\delta\bm{s}_{n,\bm{k}}\text{ ,}$ (3)
[…]
Figure 1: (a) Band structure exhibiting spin-split Dirac cones at energies $\mp J_{\text{ex}}/2$ along the out-of-plane axis $\hat{\bm{z}}$, and a band gap at charge neutrality (yellow highlight), with $J_{\text{ex}}=40\,\text{meV}$ and $\lambda_{\text{KM}}=\lambda_{\text{R}}=4\,\text{meV}$ (b) Spin texture of the valence band, showing a sharp quenching at the band edge (yellow highlight) due to Kane-Mele SOC. (c) Damping-like non-equilibrium spin texture (arbitrary units) of the valence band, showing a source of DL torque at the band edge and at the Dirac point.
In graphene this interpretation based on non-equilibrium spin dynamics is more complicated as the spin is not only coupled to a classical field, but also to the pseudospin angular momentum operator. Nevertheless, the semi-classical picture may be expanded to study the coupled spin-pseudospin dynamics given by the set of operators $\langle s_{i}\rangle$, $\langle\sigma_{j}\rangle$ and $\langle s_{i}\sigma_{j}\rangle$, with $i,j=x,y,z$; whose dynamics are determined by the Ehrenfest theorem. Strong spin-momentum locking serves as a landmark for a large DL torque, generated by the coupling between spin and an effective magnetic field. An additional DL torque landmark emerges from the coupling between spin and pseudospin, represented by spin-pseudospin entanglement hot spots at avoided band crossings. In the supplementary material we develop our methodology in detail, showing that it yields equivalent results to the Kubo formalism, while providing a deeper grasp of the subjacent microscopic processes [24].
Figure 2: $S_{\text{DL}}$ with (solid black curve) and without (dashed grey curve) Kane-Mele SOC, with $J_{\text{ex}}=40\,\text{meV}$ and $\lambda_{\text{R}}=4\,\text{meV}$. A SOT plateau is observed throughout the band gap at charge neutrality enabled by $\lambda_{\text{KM}}$ (yellow highlight). At the Dirac points (red and blue highlights) a sharp valley in $S_{\text{DL}}$ is observed due to spin-pseudospin entanglement, which can be reverted into a sharp peak by $\lambda_{\text{KM}}$. Left inset: $S_{\text{FL}}$, is insensitive to $\lambda_{\text{KM}}$. Right inset: $S_{\text{DL}}$ at the gap (yellow), spin majority Dirac point (red), and spin minority Dirac cone (blue) as a function of $\lambda_{\text{KM}}$.
The results for the non-equilibrium spin density, presented in Fig. 2, reveal that the DL spin density, $S_{\text{DL}}$, presents prominent features at the avoided band crossings, both at charge neutrality (yellow highlight), and at the spin-split Dirac points (red and blue highlights). Most remarkably, a SOT plateau is observed throughout the entire band gap at charge neutrality, where we additionally observe a quantized Hall conductivity of $\sigma_{xy}=2e^{2}/h$ corresponding to a topological quantum anomalous Hall phase, consistent with previous studies in magnetic Rashba graphene [42, 43]. Interestingly however, while the quantum anomalous Hall phase remains quantized regardless of $\lambda_{\text{KM}}$, the $S_{\text{DL}}$ plateau only forms upon breaking particle-hole symmetry by including Kane-Mele SOC, and vanishes for $\lambda_{\text{KM}}=0$ (see right inset of Fig. 2). Thus, in absence of conductive bulk states, we observe the emergence of a DL torque which must be carried by the topological edge states related to the quantum anomalous Hall regime. Both of these quantities result from a non-equilibrium modification of the Bloch states throughout the Fermi sea due to adiabatic transport in $k$-space; however, while the Hall conductivity stems from a transverse charge current [44], we show here that $S_{\text{DL}}$ derives from a non-equilibrium spin texture. This behavior reveals a SOT driven by topological physics which has been so far unnoticed.
We offer an explanation for the origin of the SOT plateau based on the semi-classical spin-pseudospin dynamics, relating it to band structure features. As shown in Fig. 1-(a), the band gap at charge neutrality (yellow highlight) is generated by the avoided crossing between the upper band of the spin majority Dirac cone and the lower band of the spin minority one, with positive and negative out-of-plane spin polarization respectively and both exhibiting the same helicity. For $\lambda_{\text{KM}}=0$ these bands present a radial pseudospin polarization along $\pm\hat{\bm{k}}$, opposite to one another, belonging to orthogonal pseudospin subspaces. Thus, the hybridization between both bands near the gap leads to a loss of their out-of-plane spin polarization due to spin-pseudospin entanglement, while their in-plane spin texture is unaffected. On the other hand, when introducing Kane-Mele SOC both bands acquire the same out-of-plane pseudospin component proportional to $\lambda_{\text{KM}}$, and thus share a common pseudospin subspace. In this case the hybridization between bands of up and down spin polarization stemming from the opposite Dirac cones modifies the in-plane spin texture [24]. This process results in a sharp variation of the spin helicity at the band edges, as highlighted in yellow in Fig. 1-(b), where the mark of strong spin-momentum locking forecasts a large $S_{\text{DL}}$. Indeed, as the Bloch states are transported under an applied electric field, those near the band edge feel a strong variation of the effective magnetic field, which drives the non-equilibrium spin dynamics resulting in a non-equilibrium spin texture source at the band edge, shown in Fig. 1-(c). This effect is opposite at the valence and conduction bands, manifesting broken particle-hole symmetry, and therefore, within the gap the fully filled valence band generates the DL SOT plateau [24].
The FL torque, on the other hand, stems from the Fermi-surface REE and remains practically insensitive to $\lambda_{\text{KM}}$, as shown in the left inset of Fig. 2. Even though the spin helicity is modified by $\lambda_{\text{KM}}$ near the gap, $S_{\text{FL}}$ remains practically unchanged as the two Fermi contours of opposite velocities cancel one another. At higher energies the inner contour shrinks until vanishing when reaching the Dirac point, where $S_{\text{FL}}$ reaches its maximum value as the band stemming from the opposite Dirac cone exclusively dominates Fermi-surface transport. Kane-Mele SOC opens a semi-gap at the Dirac point; however, $S_{\text{FL}}$ is unaffected as it is solely determined by the outer Fermi contour.
On the contrary, the DL torque shows a remarkable behavior at the Dirac points, where a large non-equilibrium spin density is generated within the Kane-Mele semi-gap (shown by the red and blue highlights in Fig. 2), along with the corresponding spin Hall effect driven by the Berry curvature [24]. Distinct to the SOT plateau at charge neutrality whose underlying mechanism relies on the enhancement of spin-momentum locking at the band edges, the spin density generated at the Dirac point stems from spin-pseudospin entanglement sources at the avoided band crossing. Rashba SOC not only generates helical spin-momentum locking, but also promotes spin-pseudospin entanglement by hybridizing the opposite Dirac cones. At $k=0$ the latter effect becomes more relevant due to the vanishing momentum, where, considering $\lambda_{\text{KM}}=0$, the bottom-most and top-most bands hybridize to form spin-pseudospin entangled states. Thus, while at a finite momentum the two bands stemming from a same Dirac cone present opposite pseudospin polarization, near $k=0$ they share a common pseudospin subspace. By these means, under an applied electric field inter-band transitions at the Dirac point generate a large non-equilibrium spin density of opposite sign to the one generated by the outer Fermi contour, resulting in sharp $S_{\text{DL}}$ valleys at both Dirac cones for $\lambda_{\text{KM}}=0$, shown by the dashed curve in Fig. 2.
The DL torque at the Dirac cones can be tuned by Kane-Mele SOC. Indeed, Rashba SOC promotes a hybridization between the spin-split Dirac cones, which are spectrally separated by $J_{\text{ex}}$ at $k=0$, imprinting an effective mass of $\approx\lambda_{\text{R}}^{2}/J_{\text{ex}}$ on both cones. On the other hand, Kane-Mele SOC induces an effective mass of opposite sign in each cone, breaking particle-hole symmetry. For $\lambda_{\text{KM}}>0$ both contributions to the effective mass collaborate at the spin minority Dirac cone, enhancing the $S_{\text{DL}}$ valley (see blue highlight in Fig.2). However, at the spin majority Dirac cone the sign of the band inversion induced by Rashba SOC is flipped by $\lambda_{\text{KM}}$. The $S_{\text{DL}}$ contribution stemming from within the Dirac point now interferes constructively with the one stemming from the outer Fermi contour, generating a pronounced peak in the non-equilibrium spin density (see red highlight in Fig.2).
In conclusion, the role of spin-pseudospin entanglement and presence of topological states have been shown to generate a novel type of spin-orbit torque driven by topological physics, and maximally optimize charge-to-spin conversion via the Rashba-Edelstein effect. We expect these conclusions to be general to systems where spin-momentum locking is mediated by additional degrees of freedom, as can also occur with orbital angular momentum [45]. Our findings open new paths for efficient spintronic devices, as well as for fundamental research on novel topological phases.
Acknowledgements.
The authors acknowledge funding from the project I+D+i PID2022-138283NB-I00 funded by MICIU/AEI/10.13039/501100011033/ and “FEDER Una manera de hacer Europa”, FLAG-ERA project MNEMOSYN (PCI2021-122035-2A) funded by MICIU/AEI /10.13039/501100011033 and the European Union NextGenerationEU/PRTR, and ERC project AI4SPIN funded by the European Union’s Horizon Europe research and innovation programme - European Research Council Executive Agency under grant agreement No 101078370-AI4SPIN. This work is also supported by MICIU with European funds‐NextGenerationEU (PRTR‐C17.I1) and by and 2021 SGR 00997, funded by Generalitat de Catalunya. J.M.D. acknowledges support from MICIU grant FPI PRE2021-097031. ICN2 is funded by the CERCA Programme/Generalitat de Catalunya and supported by the Severo Ochoa Centres of Excellence programme, Grant CEX2021-001214-S, funded by MCIN/AEI/10.13039.501100011033.